OpenStudy (anonymous):

integrate 2/(5x-1) dx

OpenStudy (anonymous):

use substitution method put 5x- 1 = t

OpenStudy (anonymous):

In uv substitution, would that be u?

OpenStudy (anonymous):

i think u r asking for integration by parts, here there is not need to apply this rule put 5x-1 = t 5dx = dt dx = dt/5 therefore (2/5t)*dt ( 2/5) ln(t) put back the value of t in final solution (2/5) ln(5x-1)

OpenStudy (anonymous):

That makes so much more sense. Thank you! So, if I wanna find whether or not this converges when you take the integral from 1 to infinity, the function diverges, right?

OpenStudy (anonymous):

as far as divergence or convergence is concerned, u need to apply leibniz integral rule........and i really dont remember the rule right now.......

OpenStudy (anonymous):

We never learned that in class.

OpenStudy (anonymous):

sorry that rule was cauchy integral test...........u must have heard of this rule

OpenStudy (anonymous):

No

OpenStudy (anonymous):

sorry i m not able to recollect

OpenStudy (anonymous):

It's okay. Thank you.